News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 5 December 2003 8.30am EST

Last week the observing schedule was halted due to high radiation associated with solar flare activity. The loads were halted on Dec 2 at 12:31pm EST through an autonomous execution of the SI Safing SCS 107 and science was resumed at 9:27am EST on Dec 4. Six science observations (total 89.6 Ks) were impacted by the event and will be rescheduled in a later load.

The IEPHIN instrument has recently shown evidence of anomalous performance during periods of elevated operating temperature. In two cases the power supply voltage dropped as it became current limited due to the high temperature then returned to its nominal value once the unit cooled following a maneuver to a new attitude. The team has been monitoring a trend towards higher temperatures in this unit for some time and work is underway to develop an alternate radiation trigger using data from the HRC anticoincidence shield, in the event IEPHIN becomes unavailable. In the short term, observations will be scheduled to ensure that the unit does not reach temperatures high enough to cause anomalous performance.

The observing schedule for the next week is shown below and includes an observation of GRS 1915+105 accepted as a Target of Opportunity on Dec 2.

------------------------------------------
Radiation Belts                      Dec 6
GRS 1915+105 (TOO)      ACIS-S       Dec 7
A267                    ACIS-I
ZCMA                    ACIS-S
ACIS Background         HRC-I        Dec 8
0540-697                ACIS-S
Radiation Belts                      Dec 9
IRAS FSC 09531+1430     ACIS-S
Her X-1                 ACIS-S/HETG
SN 1986J                ACIS-S       Dec 10
Radiation Belts                      Dec 11
Epsilon ORI             ACIS-S/HETG  Dec 12
SC1329-313              ACIS-I       Dec 13
MACS-J1226.8+2153       ACIS-I
F1525-1892              ACIS-S
GJ1245                  ACIS-S       Dec 14

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports